OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 y=-1/2x+5

OpenStudy (anonymous):

sorry. dont know this one ;/

OpenStudy (anonymous):

-2x + 4y = -7 4y = 2x - 7 y = 2/4x - 7/4 y = 1/2x - 7/4 slope here is 1/2 and y intercept is -7/4 y = -1/2x + 5 slope here is -1/2 and y intercept is 5 Learn this... if the slopes and the y intercepts are the same, then the lines will have infinite solutions if the slopes are the same and the y intercepts are different, then there is no solution if the slopes and the y intercepts are different, then there will be 1 solution so what do you think ?

OpenStudy (anonymous):

darn it, dont worry thanks for trying @kelliee

OpenStudy (anonymous):

still not sure,,,.. @kelliegirl33

OpenStudy (anonymous):

its not infinite

OpenStudy (anonymous):

you have slopes that are different and y intercepts that are different. That means there is 1 solution.

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

correct. 1 solution because the lines will intersect in exactly one spot (x,y)

OpenStudy (anonymous):

I get it now, thank you for you explanation !!

OpenStudy (anonymous):

no problem...you need to learn it from somewhere. At least just to get out of high school

OpenStudy (anonymous):

yes sigh